# Shadows and Observational Images of a Schwarzschild-like Black Hole Surrounded by a Dehnen-type Dark Matter Halo

### v1

@@ Abstract
This study investigates the optical appearance of a Schwarzschild-like black hole (BH) surrounded by a Dehnen-type dark matter (DM) halo, with a focus on how the DM halo’s density $\rho_{s}$ and radius $r_{s}$ influence the BH’s shadow and photon ring. First, the radius $r_{h}$ of the BH’s event horizon and the equation of motion for photons were derived, and observational data from the Event Horizon Telescope (EHT) for M87* were used to constrain the parameters $\rho_{s}$ and $r_{s}$ of the DM halo. Afterward, by varying the values of $\rho_{s}$ and $r_{s}$, key parameters such as the effective potential $V_{eff}$ of photons, the critical impact parameter $b_{ph}$, the radius $r_{isco}$ of the innermost stable circular orbit,and the the radius $r_{ph}$ of the photon sphere were calculated for each case. It was found that as $\rho_{s}$ and $r_{s}$ increase, the above mentioned parameters all show an increasing trend. Subsequently, we examined the optical appearance of the BH under two models: one with an optically and geometrically thin accretion disk and the other with a static spherical accretion model. The findings indicate that as $\rho_{s}$ and $r_{s}$ increase, the peak of the received intensity shifts towards a higher impact parameter $b$, resulting in a distinct optical appearance.
Keywords
shadows, optical appearance, photon ring, dark matter
preprint: APS/123-QED
I Introduction
@@
General relativity (GR) stands as one of the most successful theories in describing gravitational phenomena and has gained widespread acceptance. As one of its key theoretical predictions, BHs have remained a focal point of physics research for decades, captivating scientific attention due to their enigmatic nature. Extensive theoretical investigations have been conducted, as exemplified by studies in references Abbott et al. (2016a); Berti et al. (2009); Abbott et al. (2016b); Akiyama et al. (2019a); Banados et al. (1992); Kovtun et al. (2005); Hawking (1976, 1975); Akiyama et al. (2022); Shenker and Stanford (2014) which explore BHs from diverse perspectives. As one of the possible products of the gravitational collapse of massive stars Oppenheimer and Snyder (1939), BHs cannot be directly observed due to their intense gravitational fields, and their existence has yet to be conclusively verified. The study of BHs entered a new phase in 2019 when the Event Horizon Telescope (EHT) team unveiled the image of the supermassive BH located at the center of a galaxy Akiyama et al. (2019b, c). The image of BH M87* revealed a central dark area encircled by a luminous ring. The shaded region in the figure corresponds to the BH shadow, which appears as a dark area in the observer’s field of view after light rays fall into the BH horizon. The bright ring structure in the image includes the BH photon ring. The photon ring is formed because the gravitational field of the BH itself causes the distortion of the spacetime structure around it, causing the trajectories of photons passing near the BH to bend, thereby producing a bright ring structure in the observer’s field of view Gralla et al. (2019). This gravitational lensing effect, where light rays are deflected around the BH, not only demonstrates the profound influence of the BH on the surrounding spacetime but also provides a way to further study the properties of BHs. The release of the real images of BHs has also inspired theoretical research on BH images. As a result, a large number of studies have been carried out. For example, in references Chakhchi et al. (2022); Hioki and Maeda (2009); Nojiri and Odintsov (2025); Gallo and Mädler (2025); Yang et al. (2024a); Promsiri et al. (2024); Meng et al. (2024); Promsiri et al. (2024); Yashwanth et al. (2024); Meng et al. (2023a); Amir and Ghosh (2016); Eslam Panah and Heidari (2025); Abdujabbarov et al. (2016); Zeng et al. (2025a); Wang et al. (2025); Zeng et al. (2025b, 2023a, 2022a, 2023b); Kala et al. (2024), research on BH images under different gravitational effects has been conducted.
@@
Although GR has achieved great success in many aspects of describing gravity, it still does not include DM and dark energy. Observations of the cosmic microwave background indicate that the universe is mainly composed of $26.8\%$ DM and $68.5\%$ dark energy Aghanim et al. (2020); Hinshaw et al. (2013). Therefore, it is natural to assume that BHs do not exist completely in isolation, and there will also be DM surrounding them. Although DM has not been directly detected yet, there are already a lot of astronomical observation evidences indirectly proving its existence, such as the rotation curves of galaxy disks Rubin et al. (1980); Lelli et al. (2016), the Bullet Cluster Clowe et al. (2006); Corbelli and Salucci (2000), baryon acoustic oscillations Eisenstein et al. (2005); Komatsu et al. (2011), and so on. Therefore, the combination of DM and BHs has already become an object of great concern and high research value at present. Since DM does not participate in electromagnetic interactions Konoplya (2019), the influence of its gravitational effect on BHs cannot be obtained through direct electromagnetic observations. Therefore, one studies the optical appearance of BHs in a DM halo to explore DM’s influence on BHs. Relevant research has been conducted for this purpose Jusufi et al. (2019); Jha (2025); Macedo et al. (2024).
@@
The spacetime around a BH can be influenced by the surrounding matter field, and this influence will be reflected in the image of the BH Xu et al. (2018); Haroon et al. (2019); Guo et al. (2022); Jha (2025); Macedo et al. (2024). In addition, since BHs accrete the surrounding luminous matter, the different distributions of these accreting materials will further affect the image of the BH. However, it is difficult to simulate a realistic accretion disk theoretically. Therefore, the accretion models used in the theoretical study of BH accretion imaging are all simplified. At present, several widely adopted accretion disk models include the optically thin static accretion disk model Wang et al. (2023); Sui et al. (2024); Li and He (2021); Uniyal et al. (2023); Hu et al. (2023); Fathi and Cruz (2023); Yang et al. (2024b) and the spherical accretion model, which features a spherically symmetric distribution of the accreting matter. Li et al. (2024a); Abdujabbarov et al. (2016); Chakhchi et al. (2022); Malik et al. (2024); Hu et al. (2022); Chen and Yang (2025). The optical appearance of BHs formed by these different accretion models can, to a certain extent, reflect the spacetime geometry around the BH, providing a new window for the verification of different gravitational theories. In recent years, the imaging of BHs under various accretion models has attracted significant research attention, as exemplified by the simulation studies of shadow images under different accretion scenarios in Hu et al. (2022); Wen et al. (2023); Gao et al. (2023); Wen et al. (2023); Zeng et al. (2022b); He et al. (2022); Huang et al. (2024); Zare et al. (2024); Huang et al. (2024); Kumaran and Övgün (2023); Meng et al. (2023b); De Martino et al. (2023); Meng et al. (2024); Aslam and Saleem (2024); Wang (2023); Zeng et al. (2023c); Uniyal et al. (2024); Li et al. (2024b); Wang et al. (2024); Zare et al. (2024); Fathi and Cruz (2023).
In studies on DM modified BHs, DM can modify the shadow characteristics of BHs by influencing key parameters such as the event horizon radius $r_{h}$ and the photon sphere radius $r_{ph}$ Zeng et al. (2023b); Faraji and Rosa (2024). More specifically, in the context of optically thin and geometrically thin accretion disk models, DM induces subtle modifications in both the relative brightness and position of photon rings as demonstrated in brane-world BH models incorporating cosmological and DM parameters, where increasing the DM parameter $\beta$ enlarges both the $r_{h}$ and the $r_{ph}$, consequently affecting the shadow radius, while variations in accretion disk inclination angles with higher inclinations shrinking the luminous ring region further modify shadow and ring structures Zeng et al. (2022c). For nonspherical BHs, increasing the DM halo mass $\Delta M$ shifts photon orbits outward and significantly enlarges the shadow radius, while reducing the halo thickness $\Delta r$ (making the mass distribution more concentrated) produces an even more pronounced shadow radius expansion Faraji and Rosa (2024). Studies characterizing DM halos by their critical density $\rho_{c}$ and scale radius $R_{s}$ have revealed that increasing either parameter leads to an overall expansion of the shadow while producing minimal distortion to its shape Zeng et al. (2025a), though notably, an increase in the ideal fluid DM parameter conversely results in a reduction of the BH shadow size Atamurotov et al. (2021). Recently, a Schwarzschild-like BH solution has been proposed Al-Badawi et al. (2025), which describes an asymptotically flat BH surrounded by a Dehnen-type DM. This achievement has built a brand-new research platform, which is helpful for in-depth exploration of the various properties of BHs surrounded by DM.
@@
Therefore, the purpose of this paper is to study the optical appearance of BHs surrounded by DM under different static accretion models, and to explore the influence of DM halos on BHs through their optical appearance. Specifically, we will discuss a Schwarzschild-like solution surrounding a Dehnen-type DM halo Al-Badawi et al. (2025), and use the ray-tracing method to study the photon trajectories and photon rings around the BH. Then, we will present the optical appearance of the BH illuminated by the optically and geometrically thin static accretion disk model and the static spherical accretion model, and further discuss the influence of DM parameters on the optical appearance of the BH.
@@
The structure of this paper is as follows: In Section II , we recall the Schwarzschild-like BH surrounded by a Dehnen-type DM halo. In Section III , we calculate the equations of motion of photons along geodesics and constrain the parameters $\rho_{s}$ and $r_{s}$ of the DM halo using the observational data of M87*. In Section IV , we classify the geodesic trajectories following the method in Gralla et al. (2019), present the corresponding photon trajectory diagrams for different values of parameters $\rho_{s}$ and $r_{s}$, $r_{s}$, and study the Schwarzschild-like optical appearance of a BH surrounded by a Dehnen-type DM halo under an optically and geometrically thin accretion disk. In Section V , we study the optical image of a BH under the static spherical accretion model. Finally, in Section VI , we present the conclusions. In this work, we employ natural units where the speed of light $c$ and the gravitational constant $G$ are both set to unity $(c=G=1)$. Additionally, the black hole mass $M$ is normalized to 1 for simplicity. The metric convention (,+,+,+)(-,+,+,+)( - , + , + , + ) is adopted throughout the paper.
II A Brief Recall of Schwarzschild-like Black Hole
@@
The metric of this Schwarzschild-like BH surrounded by a Dehnen-type DM halo can be described by the following line element:
@@
Then, Fig. 2 displays the dependence of the event horizon radius on both $\rho_{s}$ and $r_{s}$ parameters. We find that the event horizon radius of the BH increases with $\rho_{s}$ and $r_{s}$, which is consistent with the behavior shown in Fig. 1.
@@
III Photon Spheres of Schwarzschild-like BH
III.1 The trajectory of photon motion and the effective potential
@@
Here, the dots denote derivatives with respect to $\tau$. In a static spherically symmetric spacetime, two conserved quantities arise due to symmetries, namely the energy and the angular momentum, which can be expressed as follows
@@
Here, the symbol “$\pm$” denotes the propagation direction of photons. The “+” corresponds to the photon moving in the clockwise direction, while the “-” represents the photon moving in the counterclockwise direction. Moreover, $b\equiv L/E$ denotes the impact parameter of the photon, and $V_{eff}$ refers to the effective potential that governs the radial motion of the photon. This can be mathematically expressed as follows:
@@
Here, $r_{ph}$ and $b_{ph}$ represent the radius of the unstable photon sphere and the critical impact parameter respectively. In Fig. 3, the corresponding curves of the effective potential varying with different values of $\rho_{s}$ and $r_{s}$ are presented.
@@
III.2 Constraints of the M87* shadow on parameters.
In the previous subsection, we discussed the photon sphere radius and the critical impact parameter of the BH, both of which vary with $\rho_{s}$ and $r_{s}$. This implies that changes in the apparent size of the BH reflect the variation of these parameters. This offers a valuable opportunity to constrain their possible ranges using observational data of BH shadows. The shadow radius of the BH, as measured by a distant observer located at infinity, is given by:
$\begin{split}r_{sh}=\frac{r_{ph}}{\sqrt{f(r_{ph})}}\end{split}.$ (14)
In 2019, the EHT released the shadow image of the supermassive BH at the center of the M87* galaxy. The results showed that the angular diameter of the shadow is $\theta_{\rm M87^{*}}=(42\pm 3)\ \mu\text{as}$ Akiyama et al. (2019c). The mass of M87* is $M_{\rm M87^{*}}=(6.5\pm 0.9)\times 10^{9}M_{\odot}$, where $M_{\odot}$ is the mass of the sun, and the distance from M87* to the Earth is $D_{\rm M87^{*}}=16.8\pm 0.8\,\text{Mpc}$. Moreover, the shadow diameter of M87* is calculated to be approximately $d_{sh}^{\rm M87^{*}}\simeq 11.0\pm 1.5$ Bambi et al. (2019). This implies that the shadow radius $r_{sh}$ should satisfy:

### v2

@@ Abstract
This paper investigates the optical appearance of a Schwarzschild-like black hole (BH) surrounded by a Dehnen-(1, 4, 5/2) type dark matter (DM) halo, with a focus on how the DM halo’s density $\rho_{s}$ and radius $r_{s}$ influence the BH’s shadow and photon ring. First, the radius $r_{h}$ of the BH’s event horizon and the equation of motion for photons were derived, and observational data from the Event Horizon Telescope (EHT) for M87* were used to constrain the parameters $\rho_{s}$ and $r_{s}$ of the DM halo. Afterward, by varying the values of $\rho_{s}$ and $r_{s}$, key parameters such as the effective potential $V_{eff}$ of photons, the critical impact parameter $b_{ph}$, the radius $r_{isco}$ of the innermost stable circular orbit, and the radius $r_{ph}$ of the photon sphere were calculated for each case. It was found that as $\rho_{s}$ and $r_{s}$ increase, the above mentioned parameters all show an increasing trend. Subsequently, we investigated the optical appearance of the BH illuminated by two types of accretion models: optically and geometrically thin disk models and spherical accretion models. The findings indicate that as $\rho_{s}$ and $r_{s}$ increase, the peak of the received intensity shifts toward a higher impact parameter $b$, resulting in a distinct optical appearance.
Keywords:
shadows; optical appearance; photon ring; dark matter
1 Introduction
@@
General relativity (GR) stands as one of the most successful theories in describing gravitational phenomena and has gained widespread acceptance. As one of its key theoretical predictions, BHs have remained a focal point of physics research for decades, captivating scientific attention due to their enigmatic nature. Extensive theoretical investigations have been conducted, as exemplified by studies in references LIGOScientific:2016aoc ; Berti:2009kk ; LIGOScientific:2016lio ; EventHorizonTelescope:2019dse ; Banados:1992wn ; Kovtun:2004de ; Hawking:1976de ; Hawking:1975vcx ; EventHorizonTelescope:2022wkp ; Shenker:2013pqa which explore BHs from diverse perspectives. As one of the possible products of the gravitational collapse of massive stars Oppenheimer:1939ue , BHs cannot be directly observed due to their extremely strong gravitational fields, and thus can only be studied through indirect approaches. The study of BHs entered a new phase in 2019 when the EHT team unveiled the image of the supermassive BH located at the center of a galaxy EventHorizonTelescope:2019uob ; EventHorizonTelescope:2019ggy . The image of BH M87* revealed a central dark area encircled by a luminous ring. The shaded region in the figure corresponds to the BH shadow, which appears as a dark area in the observer’s field of view after light rays fall into the BH horizon. The bright ring structure in the image includes the BH photon ring. The photon ring is formed because the gravitational field of the BH itself causes the distortion of the spacetime structure around it, causing the trajectories of photons passing near the BH to bend, thereby producing a bright ring structure in the observer’s field of view Gralla:2019xty . This gravitational lensing effect, where light rays are deflected around the BH, not only demonstrates the profound influence of the BH on the surrounding spacetime but also provides a way to further study the properties of BHs. The release of real BH images has not only inspired theoretical research on BH imaging but also led to extensive subsequent studies. Subsequently, comprehensive investigations of BH images under various gravitational effects have been conducted Amir:2016cen ; Hioki:2009na ; Abdujabbarov:2016hnw ; Chakhchi:2022fls ; Zeng:2022fdm ; Zeng:2022pvb ; Yang:2023tip ; Promsiri:2023rez ; Meng:2023uws ; Zeng:2023zlf ; EslamPanah:2024gxx ; Gallo:2024wju ; Kala:2024fvg ; Meng:2024puu ; Nojiri:2024qgx ; Zeng:2024ptv ; Wang:2025ihg ; Zeng:2025kqw ; Yashwanth:2024suw .
@@
Although GR has achieved great success in many aspects of describing gravity, it still does not include DM and dark energy. Observations of the cosmic microwave background indicate that the universe is mainly composed of $26.8\%$ DM and $68.5\%$ dark energy Planck:2018vyg ; WMAP:2012nax . Therefore, it is natural to assume that BHs do not exist completely in isolation, and there will also be DM surrounding them. Although DM has not been directly detected yet, there are already a lot of astronomical observation evidences indirectly proving its existence, such as the rotation curves of galaxy disks Rubin:1980zd ; Lelli:2016zqa , the Bullet Cluster Clowe:2006eq ; Corbelli:1999af , baryon acoustic oscillations SDSS:2005xqv ; WMAP:2010qai , and so on. Therefore, the combination of DM and BHs has already become an object of great concern and high research value at present. In the study of interactions between DM and BHs, various analytical models have been proposed to describe DM BH solutions Graham:2005xx ; Dutton:2014xda ; Navarro:1995iw ; Burkert:1995yz ; Dehnen:1993uh ; Gohain:2024eer ; Pantig:2022whj . One widely used model is the Dehnen model Dehnen:1993uh . Proposed by Dehnen in 1993, this model features a spherically symmetric density profile determined by three parameters ($\alpha$, $\beta$, $\gamma$), designed to simulate the structure and dynamics of elliptical galaxies and galactic bulges Dehnen:1993uh . The gravitational potential corresponding to the Dehnen model is analytical for all values of $\gamma$ . In the specific case where $\gamma=3/2$, the Dehnen model most closely matches de Vaucouleurs’ $R^{1/4}$ empirical profile—used to describe the surface brightness of elliptical galaxies and bulges—in terms of surface density and distribution function Dehnen:1993uh . Consequently, the Dehnen model has been widely extended to construct DM halos surrounding BHs. Extensive studies have been conducted on the properties of BH solutions surrounded by Dehnen-type DM halos, including thermodynamic characteristics and quasinormal modes Alloqulov:2025ucf ; Al-Badawi:2025njy ; Al-Badawi:2024qpt ; Jha:2024ltc . Moreover, DM does not participate in electromagnetic interactions in most scenarios Konoplya:2019sns . As a result, the gravitational influence of DM on BHs cannot be directly probed through electromagnetic observations. Therefore, one studies the optical appearance of BHs in a DM halo to explore DM’s influence on BHs. Relevant research has been conducted for this purpose Jusufi:2019nrn ; Jha:2024ltc ; Macedo:2024qky ; Faraji:2024ein ; Zeng:2021mok ; Atamurotov:2021hoq .
@@
The spacetime around a BH can be influenced by the surrounding matter field, and this influence will be reflected in the image of the BH Xu:2018wow ; Haroon:2018ryd ; Guo:2022rql ; Jha:2024ltc ; Macedo:2024qky . In addition, since BHs accrete the surrounding luminous matter, the different distributions of these accreting materials will further affect the image of the BH. However, it is difficult to simulate a realistic accretion disk theoretically. Therefore, the accretion models used in the theoretical study of BH accretion imaging are all simplified. At present, several widely adopted accretion disk models include the optically thin static accretion disk model Wang:2023vcv ; Sui:2023tje ; Li:2021riw ; Uniyal:2022vdu ; Hu:2023bzy ; Fathi:2023ccx ; Yang:2024utv and the spherical accretion model, which features a spherically symmetric distribution of the accreting matter. Li:2024abk ; Abdujabbarov:2016hnw ; Chakhchi:2022fls ; Malik:2024foi ; Hu:2022lek ; Chen:2025ifv . The optical appearance of BHs formed by these different accretion models can, to a certain extent, reflect the spacetime geometry around the BH, providing a new window for the verification of different gravitational theories. In recent years, the imaging of BHs under various accretion models has attracted significant research attention, as exemplified by the simulation studies of shadow images under different accretion scenarios in Hu:2022lek ; Wen:2022hkv ; Gao:2023mjb ; Zeng:2021dlj ; He:2021htq ; Huang:2024bbs ; Zare:2024dtf ; Kumaran:2023brp ; Meng:2023htc ; DeMartino:2023ovj ; Meng:2024puu ; Aslam:2024mhn ; Wang:2023rjl ; Zeng:2023fqy ; Uniyal:2023ahv ; Li:2024owp ; Wang:2024lte ; Fathi:2023ccx .
Recently, Ahmad Al-Badawi et al. proposed a novel Schwarzschild-like solution describing an asymptotically flat Schwarzschild BH surrounded by Dehnen-(1, 4, 5/2) type DM Al-Badawi:2024asn . After analyzing the curvature properties and energy conditions of this solution to investigate the BH characteristics, they conducted a systematic study of the timelike geodesics around the BH, laying the foundation for subsequent research on the imaging features of accretion disks around such BHs. Follow-up studies have systematically examined other properties of this BH Liang:2025vux ; Hamil:2025pte ; Rani:2025esb . Among them, Rani et al. Rani:2025esb employed the ray-tracing method to investigate the optical appearance of the BH illuminated by an optically thin accretion disk. However, their work only considered a specific thin-disk model (with a radiation profile following a Gaussian function) and did not incorporate existing observational data to constrain the DM parameters. This motivates us to conduct a more comprehensive study on the optical appearance of this BH solution under different static accretion disk and spherical accretion models, while constraining the DM parameters with available observational data. By providing optical images under various accretion scenarios, we aim to extend the understanding of the observational signatures of this BH model.
@@
Therefore, the purpose of this paper is to investigate the optical appearance of a Schwarzschild-like BH immersed in Dehnen-type DM under both static thin disk accretion models and spherical accretion scenarios, and to explore the influence of DM halos on BHs through these optical signatures. Specifically, we first examine the photon sphere around a Schwarzschild-like BH solution surrounded by a Dehnen-type DM halo Al-Badawi:2024asn . Using the ray-tracing method, we investigate the photon trajectories around the BH and constrain the DM parameters with existing EHT observational data of M87*. Within these constrained parameter ranges, we classify the photon trajectories around this Schwarzschild-like BH. Finally, we analyze the optical appearance under both optically and geometrically thin accretion disks and spherical accretion scenarios.
@@
The structure of this paper is as follows: In Section 2 , we recall the Schwarzschild-like BH surrounded by a Dehnen-type DM halo. In Section 3 , we calculate the equations of motion of photons along geodesics and constrain the parameters $\rho_{s}$ and $r_{s}$ of the DM halo using the observational data of M87*. In Section 4 , we classify the geodesic trajectories following the method in Gralla:2019xty , present the corresponding photon trajectory diagrams for different values of parameters $\rho_{s}$ and $r_{s}$, and study the Schwarzschild-like optical appearance of a BH surrounded by a Dehnen-type DM halo under an optically and geometrically thin accretion disk. In Section 5, we study the optical images of a BH under thin spherical accretion models. Finally, in Section 6 , we present the conclusions. In this work, we employ natural units where the speed of light $c$ and the gravitational constant $G$ are both set to unity $(c=G=1)$. Additionally, the BH mass $M$ is set to 1 for computational purposes. The metric convention (,+,+,+)(-,+,+,+) is adopted throughout the paper.
2 A brief recall of Schwarzschild-like black hole
@@
The metric of this Schwarzschild-like BH surrounded by a Dehnen-(1, 4, 5/2) type DM halo can be described by the following line element:
@@
Then, Fig. 2 displays the dependence of the event horizon radius on both $\rho_{s}$ and $r_{s}$ parameters. The metric functions and the variation of the horizon radius with the DM parameters are shown in Figures 1 and 2, respectively, which are consistent with the results reported in Al-Badawi:2024asn . Therefore, we will not discuss them in further detail here.
@@
3 Photon spheres of Schwarzschild-like black hole
3.1 The trajectory of photon motion and the effective potential
@@
Here, the dots denote derivatives with respect to $\tau$. In a static spherically symmetric spacetime, two conserved quantities arise due to symmetries, namely the energy $E$ and the angular momentum $L$, which can be expressed as follows
@@
Here, the symbol $\pm$” denotes the propagation direction of photons. The “++” corresponds to the photon moving in the clockwise direction, while the “-” represents the photon moving in the counterclockwise direction. Moreover, $b\equiv L/E$ denotes the impact parameter of the photon, and $V_{eff}$ refers to the effective potential that governs the radial motion of the photon. This can be mathematically expressed as follows:
@@
Here, $r_{ph}$ and $b_{ph}$ represent the radius of the unstable photon sphere and the critical impact parameter respectively. And the prime notation denotes derivative with respect to $r$. Hence, we obtain
$\begin{split}b_{ph}=\frac{r_{ph}}{\sqrt{f(r_{ph})}}.\end{split}$ (14)
In Fig. 3, the corresponding curves of the effective potential varying with different values of $\rho_{s}$ and $r_{s}$ are presented.
@@
3.2 Constraints of the M87* shadow on parameters
In the previous subsection, we discussed the photon sphere radius $r_{ph}$ and the critical impact parameter $b_{ph}$ of the BH, both of which vary with $\rho_{s}$ and $r_{s}$. The variation of $b_{ph}$ reflects changes in the apparent size of the BH in the observer’s frame, thereby offering a valuable opportunity to constrain the possible parameter ranges through observational data of BH shadows. The shadow radius $r_{sh}$ of the BH, as measured by a distant observer located at infinity, is given by
$\begin{split}r_{sh}\equiv b_{ph}=\frac{r_{ph}}{\sqrt{f(r_{ph})}}\end{split}.$ (15)
In 2019, the EHT released the shadow image of the supermassive BH at the center of the M87* galaxy. The results showed that the angular diameter of the shadow is $\theta_{\rm M87^{*}}=(42\pm 3)\ \mu\text{as}$ EventHorizonTelescope:2019ggy . The mass of M87* is $M_{\rm M87^{*}}=(6.5\pm 0.9)\times 10^{9}M_{\odot}$, where $M_{\odot}$ is the mass of the sun, and the distance from M87* to the Earth is $D_{\rm M87^{*}}=16.8\pm 0.8\,\text{Mpc}$. Moreover, the shadow diameter of M87* is calculated to be approximately $d_{sh}^{\rm M87^{*}}\simeq 11.0\pm 1.5$ Bambi:2019tjh . This implies that the shadow radius $r_{sh}$ should satisfy: